### v1

# Graphical Calculus for Non-Gaussian Quantum States

Abstract
We provide a graphical method to describe and analyze non-Gaussian quantum states using a hypergraph framework. These states are pivotal resources for quantum computing, communication, and metrology, but their characterization is hindered by their complex high-order correlations. The formalism encapsulates transformation rules for any Gaussian unitary operation and local quadrature measurement, offering a visually intuitive tool for manipulating such states through experimentally feasible pathways. Notably, we develop methods for the generation of complex hypergraph states with more or higher-order hyperedges from simple structures through Gaussian operations only, facilitated by our graphical rules. We present illustrative examples on the preparation of non-Gaussian states rooted in these graph-based formalisms, revealing their potential to advance continuous-variable general quantum computing capabilities.
Introduction.— In the quest to demonstrate a genuine quantum advantage in the continuous variable (CV) domain, substantial progress on the controllable generation of multimode non-Gaussian states has recently been made by photon subtraction  [1, 2, 3], nonlinear unitary operations, such as spontaneous parametric down- conversion [4] or Kerr interactions in microwave cavities coupled superconducting artificial atoms [5, 6]. Such strategies usually introduce higher-order statistical moments of the quadrature field operators, go beyond Gaussian states and therefore have irreplaceable advantages as quantum resource [7, 8] in various tasks, such as entanglement distillation [9, 10, 11], quantum error correction [12, 13], quantum metrology [6, 14], and CV quantum computation [15, 16, 17]. Compared to the remarkable progress in experimental realizations, how to understand and more definitely classify the multimode non-Gaussian states with high-order interplay among quadratures, still faces a formidable challenge. To address it in a more intuitive and systematic manner, the adoption of graphical formalisms [18, 19, 20, 21, 22, 23, 24] as a means of describing quantum states emerges as a highly promising solution.
Figure 1: Examples of graph and hypergraph states, as well as the graphical calculus for hypergraph states as discussed in this paper. (a) Topology of a four-mode square graph state and the circuit to generate it. (b) A four-mode hypergraph state. Hypergraph states generalize graph states by allowing hyperedges (colored closed shapes), which represent interactions between multiple modes. (c) Based on the graphical rules proposed in this paper, a complex hypergraph state with more or higher-order hyperedges can be generated from simple structures solely through Gaussian operations.
Graph states [25, 26, 27] and hypergraph states [28, 29, 30, 31, 32] were first defined in discrete-variable (DV) systems to form multiparticle entangled states corresponding to mathematical graphs. They have applications in various contexts, ranging from quantum error correction [33, 34] to measurement-based quantum computation [35, 36]. This kind of definition in terms of graphical formalism then got generalised to CV system [37, 38], and examples are shown in Fig. 1(a) and Fig. 1(b). It was proved that any Gaussian state can be described by a CV graph state [39], as the first and second-order statistical moments of the quadratures which completely characterise a Gaussian state can be derived from the edges in the corresponding graph. Naturally, CV hypergraph states have richer graphical structures [40, 41], since they include multimode interactions. Such inherent multimode nonlinearity can be used to represent the high-order correlation in non-Gaussian states. Ref. [40] showed an evidence that a hypergraph state restricted to a Gaussian measurement strategy can be a candidate for universal quantum computing. Nevertheless, the ramification of arbitrary Gaussian operations on non-Gaussian hypergraph states remains an underexplored frontier, necessitating further investigation.
In this paper, we provide a unified graphical calculus that serves as a comprehensive framework for manipulating non-Gaussian states in CV systems. This calculus encapsulates hypergraph transformation rules encompassing all single-mode and typical two-mode Gaussian unitary operations, together with local quadrature measurements, offering an intuitive visual modality for implementing experimentally feasible operations on CV non-Gaussian states. Crucially, as shown in Fig. 1(c), we demonstrate the capability to generate complex hypergraph states with more or higher-order hyperedges from simple structures solely through Gaussian operations, underscoring the power and versatility of the graphical rules provided here. Furthermore, we give three examples of non-Gaussian state preparation schemes rooted in our graph-based formalism, which hold promise for application in the broader landscape of CV universal quantum computing. These instances illustrate the practical potential of the graphical calculus, positioning it as a valuable tool for advancing the frontiers of quantum information processing.
CV Hypergraph States.— A weighted hypergraph is a tuple $H=(V,E,T)$ with a set $V=\{v_{i}\}_{i=1}^{n}$ of $n$ vertices, a set $E=\{e_{j}\}_{j=1}^{|E|}$ of ${|E|}$ hyperedges, and a set $T=\{t_{e_{j}}\}_{j=1}^{|E|}$ of real valued weights associated to hyperedges, where $|E|$ denotes the number of elements of $E$. A CV hypergraph state |$\ket{H}$ is defined as an $n$-mode quantum state where the vertices represent the modes of the state and the weighted hyperedges represent generalized controlled-$Z$ gates performed on the modes [41, 40]. More formally, we have |$\ket{H}=\prod_{e_{j}\in E}C_{e_{j}}(t_{e_{j}})\ket{0}_{p}^{\otimes n}$, where |$\ket{0}_{p}$ is the zero-momentum eigenstate in the limit of infinite squeezing with normalization $\bra{p^{\prime}}\ket{p}=\delta_{p,p^{\prime}}$ and $C_{e_{j}}(t_{e_{j}})=\operatorname{e}^{it_{e_{j}}\operatorname{\mathit{\hat{q}}}_{v_{1}}\operatorname{\mathit{\hat{q}}}_{v_{2}}\dots\operatorname{\mathit{\hat{q}}}_{v_{m}}}$ is a generalized controlled-$Z$ gate acting on $m$ modes connected by the $j-$th hyperedge $e_{j}=\{v_{1},v_{2},\dots v_{m}\}$. If a hyperedge only contains two vertices, i.e. $\lvert e\rvert=2$, $C_{e}$ is the standard controlled-$Z$ gate $C_{e}(t)=\operatorname{e}^{it\operatorname{\mathit{\hat{q}}}_{v_{1}}\operatorname{\mathit{\hat{q}}}_{v_{2}}}$ and if it contains three vertices, it is the Toffoli gate $C_{e}(t)=\operatorname{e}^{it\operatorname{\mathit{\hat{q}}}_{v_{1}}\operatorname{\mathit{\hat{q}}}_{v_{2}}\operatorname{\mathit{\hat{q}}}_{v_{3}}}$ [42]. We use the shorthand notation $q_{e_{j}}\coloneqq\operatorname{\mathit{\hat{q}}}_{v_{1}}\operatorname{\mathit{\hat{q}}}_{v_{2}}\dots\operatorname{\mathit{\hat{q}}}_{v_{m}}$. A hypergraph state of $k$-th order is classified by the maximum number of $k$ modes connected with one hyperedge.
For clarity, we give an example of a 4-mode hypergraph state containing $|E|=2$ hyperedges, as shown in Fig. 2(a). One hyperedge $e_{1}=\{A,B,C\}$ has weight $t_{e_{1}}=1$ and the other hyperedge $e_{2}=\{A,D\}$ has weight $t_{e_{2}}=2$. This hypergraph state is given by
|$\ket{H}=\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}}\operatorname{e}^{i2\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{q}}}_{D}}\ket{0}_{p}^{\otimes 4}.$ (1)
Figure 2: Examples of a weighted hypergraph state and different Gaussian operations applied on it. Roman labels represent the weight of each hyperedge. The orange circle represents the initial hyperedge, and the green circle represents the newly created hyperedge. The purple solid lines and circles represent the 2nd-order and 1st-order hyperedges respectively. The dotted circles represent the higher-order moments. (a) A hypergraph $H$, corresponding to four modes and two hyperedges: one 2nd-order hyperedge $\{A,D\}$ with weight 2 and the other 3rd-order hyperedge $\{A,B,C\}$ with weight 1. On the corresponding quantum state, one can apply different Gaussian operations: (b) Applying momentum displacement $Z^{B}(-2)$ on mode $B$ adds an hyperedge representing $\operatorname{e}^{-i2\operatorname{\mathit{\hat{q}}}_{B}}$ on it. (c) Applying position displacement $X^{A}(1)$ on mode $A$ generates new hyperedges: a 1st-order hyperedge representing the term $\operatorname{e}^{-i2\operatorname{\mathit{\hat{q}}}_{D}}$ on $D$ and a 2nd-order hyperedge representing $\operatorname{e}^{-i\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}}$ for $\{B,C\}$. (d) Applying position shearing on mode $B$, $D_{q}^{B}(2)$, adds a decoration representing $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{B}^{2}}$ on $B$, which is not part of the formalism. (e) Similarly, momentum shearing $D_{p}^{A}(1)$ on mode $A$ adds decorations, representing $\operatorname{e}^{i2\operatorname{\mathit{\hat{q}}}_{D}^{2}}$ and $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{B}^{2}\operatorname{\mathit{\hat{q}}}_{C}^{2}/2}$, which fall outside the formalism on $D$ and $B,C$, but also induces a valid new hyperedge representing $\operatorname{e}^{i2\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}\operatorname{\mathit{\hat{q}}}_{D}}$ for $\{B,C,D\}$. (f) Squeezing $S^{A}(1)$ on mode $A$ changes the weights on all hyperedges adjacent to mode $A$. (g) Rotation $R^{D}(\pi)$ on mode $D$ changes the weights by inverting all signs on hyperedges adjacent to $D$. (h) Controlled-$Z$ gate $C_{A,D}(0.5)$ on mode $A$ and mode $D$ adjusts their weight by 0.5.
In order to present the graphical calculus more conveniently, we first introduce some basic rules of set operations. The hyperedge set $E_{a}$ is defined as the set of all hyperedges containing mode $a$: $E_{a}=\{e\mid e\in E,a\in e\}$. Furthermore, the adjacency set $\operatorname{\mathcal{A}}(a)$ of vertex $a\in V$ is defined as $\operatorname{\mathcal{A}}(a)=\{e\setminus\{a\}\mid e\in E_{a}\}$. To consider the bijective mapping between hyperedge and weight sets, we define the union $(E_{U};T_{U})=(E_{1};T_{1})\cup(E_{2};T_{2})$ as a union of all tuples. For hyperedges appearing in both sets, we add up their weights. More precisely, for any $(e,t_{1})\in(E_{1};T_{1})$ and $(e,t_{2})\in(E_{2};T_{2})$, we get $(e,t_{1}+t_{2})\in(E_{U};T_{U})$. For the hypergraph in Fig. 2(a), the set of hyperedges containing mode $A$ is given by $E_{A}=\{\{A,B,C\},\{A,D\}\}$ with associated weights $T_{A}=\{1,2\}$ and $\mathcal{A}(A)=\{\{B,C\},\{D\}\}$. If we want to unite $E_{A}$ and $E_{B}=\{\{A,B,C\}\}$ with $T_{B}=\{1\}$, we obtain $(E_{U};T_{U})=(\{\{A,B,C\},\{A,D\}\};\{2,2\}).$
Although not all quantum states can be transformed under the graphical rules, some states are able to be represented indirectly in the form of hypergraphs, for example the cubic phase state $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}^{3}}\ket{0}_{p}$ which contains a third-order moment term of one single-mode operator and plays important roles in universal quantum computing [43, 15]. It turns out that it can be prepared in our formalism.
Gaussian Operations on CV Hypergraph States.— In this part we provide unified graphical rules of Gaussian unitary operations and Gaussian measurements for CV hypergraph states. We show the graphical rules of single-mode Gaussian operations including displacement, shearing, squeezing, and rotation [44, 45] and further discuss the CV controlled-$Z$ gate. Derivations are given in the appendix. As any multimode Gaussian operation can be represented as a concatenation of a beam splitter, single-mode squeezing, and phase shifts through the Bloch-Messiah decomposition [46], and the beam-splitter operation can be replaced by controlled-$Z$ gate operations [42], any multimode Gaussian operation can be represented in our graphical calculus.
Operation I: Displacement.— The displacement operators include the momentum displacement $Z^{a}(s)=\operatorname{e}^{is\operatorname{\mathit{\hat{q}}}_{a}}$ and the position displacement $X^{a}(s)=\operatorname{e}^{-is\operatorname{\mathit{\hat{p}}}_{a}}$, where $s$ represents the distance that the operator $\operatorname{\mathit{\hat{q}}}_{a}$ or $\operatorname{\mathit{\hat{p}}}_{a}$ is moved in phase space [47]. Thus, the momentum displacement operator $Z^{a}(s)=\operatorname{e}^{is\operatorname{\mathit{\hat{q}}}_{a}}$ on mode $a$ modifies the hyperedge set by $(E^{\prime};T^{\prime})=(E;T)\cup(\{a\};s)$. In other words, if mode $a$ has its own hyperedge initially, this operation will increase the weight by $s$. If not, a new hyperedge with the weight $s$ will be added to the system. Fig. 2(b) shows an example of performing operation $Z^{B}(-2)$ on the hypergraph state depicted in Fig. 2(a).
However, when performing the position displacement $X^{a}(s)=\operatorname{e}^{-is\operatorname{\mathit{\hat{p}}}_{a}}$ to the hypergraph state, due to the incompatibility between two operators $\operatorname{\mathit{\hat{q}}}_{a}$ and $\operatorname{\mathit{\hat{p}}}_{a}$, additional correlation terms will be generated when they simultaneously act on the zero-momentum eigenstate |$\ket{0}_{p}$. It influences hyperedges of the adjacency $\mathcal{A}(a)$ of mode $a$ such that $(E^{\prime},T^{\prime})=(E,T)\cup(\operatorname{\mathcal{A}}(a),\{\tilde{t}\})$, where $\tilde{t}\coloneqq-s\times t_{e\cup a}$, which means the position displacement on mode $a$ will add or change hyperedges between the adjacent modes to mode $a$ with the weight multiplied by $-s$. For example, when the position displacement operator $X^{A}(1)$ is applied to mode $A$ of the hypergraph state corresponding to Fig. 2(a), two new hyperedges between vertices $\{B,C\}$ and $\{D\}$ of $\operatorname{\mathcal{A}}(a)$ are created, as represented by the purple line and purple circle in Fig. 2(c).
From the above analysis, it can be seen that in a hypergraph state of $k$-th order, displacement operations can create new hyperedges of at most order $k-1$, which is similar to rules holding for multi-qubit hypergraph states [31, 48]. Note that for CV graph states, the order of edges is $k=2$ and therefore displacement operations only cause local effects in a certain mode.
Operation II: Shearing.— The shearing operator includes the momentum shearing $D_{p}^{a}(s)=\operatorname{e}^{i\frac{s}{2}\operatorname{\mathit{\hat{p}}}_{a}^{2}}$ and the position shearing $D_{q}^{a}(s)=\operatorname{e}^{i\frac{s}{2}\operatorname{\mathit{\hat{q}}}_{a}^{2}}$, which shears the state with respect to the $\operatorname{\mathit{\hat{p}}}$ or $\operatorname{\mathit{\hat{q}}}$-axis by a gradient of $s$. The shearing operator introduces an additional second-order moment term in the exponent, which does not occur in a standard hypergraph state, so this operation leaves the set of hypergraph states. But this is in line with observations made on hypergraph states with qubits [48]. For the position shearing operation, a term without effecting the hyperedge set of the state, $D_{q}^{a}(s)\ket{H}=\operatorname{e}^{i\frac{s}{2}\operatorname{\mathit{\hat{q}}}_{a}^{2}}\ket{H}$, will be created. For example, performing the position shearing $D_{q}^{B}(2)$ on mode $B$ will change the hypergraph into the one shown in Fig. 2(d). The dotted line indicates that there is a second-order moment on its own.
Similarly, due to the incompatibility between two operators $\operatorname{\mathit{\hat{q}}}_{a}$ and $\operatorname{\mathit{\hat{p}}}_{a}$, applying the shearing in momentum operator $D_{p}^{a}(s)$ on mode $a$ will introduce new terms acting on the adjacency $\operatorname{\mathcal{A}}(a)$. More precisely, we have
$D_{p}^{a}$ $(s)\ket{H}=e^{i\frac{s}{2}(\sum_{e\in\operatorname{\mathcal{A}}(a)}t_{e\cup a}\operatorname{\mathit{\hat{q}}}_{e})^{2}}\ket{H}.$ (2)
Momentum shearing operation leads to a state which is not a hypergraph state in most cases. In the example of Fig. 2(a), the adjacency of $A$ is $\{\{B,C\},\{D\}\}$. The operation $D_{p}^{A}(1)$ on mode $A$ introduces the term $\operatorname{e}^{i(\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}+2\operatorname{\mathit{\hat{q}}}_{D})^{2}/2}$ to the original hypergraph state. Expanding the square term we get a forth-order moment term $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{B}^{2}\operatorname{\mathit{\hat{q}}}_{C}^{2}/2}$, a second-order moment term $\operatorname{e}^{i2\operatorname{\mathit{\hat{q}}}_{D}^{2}}$, and a hyperedge $\operatorname{e}^{i2\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}\operatorname{\mathit{\hat{q}}}_{D}}$ among modes $B$, $C$ and $D$, as shown in Fig. 2(e).
Therefore, it can be seen that in a hypergraph state of $k$-th order, by shearing operations, a new hyperedge of $k$-th order can be generated to connect more different modes. This effect can also link more pairs in the graph state. In principle, if a hypergraph state contains one $k$-th order hyperedge and another $l$-th order hyperedge which intersect in exactly one mode, shearing operations acting on this shared mode can expand the maximum order to $(k+l-2)$-th, offering a practical approach to generate more or higher-order hypergraph states with multimode nonlinearity using only local Gaussian operations.
Operation III: Squeezing.— Applying a one-mode squeezing operator $S^{a}(s)=\operatorname{e}^{-i\frac{s}{2}(\operatorname{\mathit{\hat{p}}}_{a}\operatorname{\mathit{\hat{q}}}_{a}+\operatorname{\mathit{\hat{q}}}_{a}\operatorname{\mathit{\hat{p}}}_{a})}$ on mode $a$ multiplies the weight of all hyperedges linked to mode $a$ by an additional factor of $\operatorname{e}^{-s}$, while other hyperedges remain unchanged. In Fig. 2(f), we show an example by squeezing mode $A$ with $S^{A}(1)$, hence, $t_{AD}=2\operatorname{e}^{-1},t_{ABC}=\operatorname{e}^{-1}$. Therefore, by using a squeezing operation, the value of weights in the hyperedges can be effectively proportionally adjusted, that is, the strength of the correlation between different modes can be changed.
Operation IV: Rotations.— Applying a local rotation $R_{a}(s)=\operatorname{e}^{i\frac{s}{2}(\operatorname{\mathit{\hat{q}}}_{a}^{2}+\operatorname{\mathit{\hat{p}}}_{a}^{2})}$ in phase space usually results in the final state no longer being a standard hypergraph state. But in some special cases, $s=n\pi$, we can still get a hypergraph state. However, the hypergraph keeps unchanged for $n$ is even, while the weights of hyperedges in $E_{a}$ will be negated for $n$ is odd. In Fig. 2(g), we show the effect of rotation $R^{D}(\pi)$ on vertex(mode) $D$. This is also consistent with the observations in the case of hypergraph states with qubits [49].
For the general case $s\neq n\pi$, the specific form of the hypergraph state after rotation operation is given in the appendix. Since graph and hypergraph states originate from a series of logic gates acting on the zero-momentum eigenstates, but rotation operations may cause a shift to the zero-position eigenstate, some complicated correlation terms will be introduced by such a shift. Therefore, the use of rotation operations in the context of manipulating hypergraph states is not straightforward.
Operation V: Two-mode Gaussian operations.— We have previously described the graphical effects of single-mode Gaussian operations on the structure of the hypergraph. It has been proven that any two-mode Gaussian operation can be decomposed into a CV controlled-$Z$ gate and a series of single-mode operations [46, 42]. So it suffices to focus on this two-mode Gaussian operation. A controlled-$Z$ gate $C_{i,j}(s)=\operatorname{e}^{is\operatorname{\mathit{\hat{q}}}_{i}\operatorname{\mathit{\hat{q}}}_{j}}$ can connect two modes $i$ and $j$ by an edge of strength $s$ or change its weight if edge $\{i,j\}$ was there before. In Fig. 2 (h), we show the effect of $C_{A,D}(0.5)$ on vertexes(modes) $A$ and $D$.
Therefore, the controlled-$Z$ gate provides a way to adjust the adjacency set of any vertex. By combining it with single-mode operations, we can create or delete hyperedges by Gaussian operation.
Operation VI: Gaussian measurements.— Since a mode collapses after measurement, measuring |$\outerproduct{m}{m}_{q,a}$ will disconnect mode $a$ from the hypergraph and delete mode $a$ from all hyperedges which were adjacent to mode $a$ before. The hyperedge set and weight set change to $(E^{\prime};T^{\prime})=(E\setminus E_{a};T)\cup(\operatorname{\mathcal{A}}(a);\{\bar{t}\})$, where the weights of all hyperedges in $\operatorname{\mathcal{A}}(a)$ is $\bar{t}\coloneqq m\times t_{e\cup a}$.
Measuring |$\outerproduct{m}{m}_{p,a}$ disconnects mode $a$ and modifies hyperedges which were previously adjacent to mode $a$. As for $q$-measurements, we remove mode $a$ from hyperedges in $E_{a}$ but then average over all possible weights. Hence, the generating state is given by
[…]
$\prod_{e^{\prime}\in E_{\neg a}}C_{e^{\prime}}(t_{e^{\prime}})\ket{0}_{p}^{\otimes\lvert V\rvert-1},$ (3)
where $E_{\neg a}=\{e\mid e\in E,a\notin e\}$ is the set of hyperedges without mode $a$.
Figure 3: Generation of a hypergraph state with two 3rd-order hyperedges starting from a hypergraph state with only one 3rd-order hyperedge by Gaussian operations. (a) The initial hypergraph state |$\ket{H}$. (b) After performing momentum shearing on mode $A$, the desired green hyperedge of $\{B,C,D\}$ is generated. Additionally, two purple dashed lines, $\operatorname{e}^{i2\operatorname{\mathit{\hat{q}}}_{D}^{2}}$ and $e^{i\operatorname{\mathit{\hat{q}}}_{B}^{2}\operatorname{\mathit{\hat{q}}}_{C}^{2}/2}$, are produced as by-products. (c) Removing the $\{A,D\}$ edge by performing a two-mode controlled-$Z$ gate $C_{A,D}(-2)$. (d) After performing the shearing operations towards the opposite direction on modes $A$ and $D$ respectively, the new hypergraph state $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}}\operatorname{e}^{i2\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}\operatorname{\mathit{\hat{q}}}_{D}}\ket{0}_{p}^{\otimes 4}$ is obtained.
Experimental feasibility and applications.— Based on the graphical calculus, it is possible to generate hypergraph states with richer structures from an initial 3rd-order hypergraph state using Gaussian operations. To realize the 3rd-order hypergraph state $e^{ig\tau\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}}\ket{0}_{p}^{\otimes 3}$ we need to construct a Hamiltonian $H=g\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}$, and act on three zero-momentum eigenstates for the evolution time $\tau$. Recently, a three-mode non-Gaussian state has been experimentally achieved via three-photon spontaneous parametric down-conversion starting from a vacuum state in superconducting systems [4]. When the system evolves from the squeezed state, the obtained state will approximate a hypergraph state as the squeezing level increases.
The ability to generate non-Gaussian states with higher multimode nonlinearity is seen as a key resource in many applications of quantum theory. Directly executing non-Gaussian operations across multiple modes proves to be highly challenging in practice. In Fig. 3, we show how to generate an additional 3rd-order hyperedge among modes $\{B,C,D\}$ from the initial hypergraph state as shown in Fig. 3(a). That is, performing a series of Gaussian operations to implement an equivalent result of a typical non-Gaussian operation, CV Toffoli gate, $C_{B,C,D}(2)=\operatorname{e}^{i2\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}\operatorname{\mathit{\hat{q}}}_{D}}$, among them [42]. More explicitly, we first use a local shearing operation on the position operator of mode $A$ to create a new 3rd-order hyperedge among its adjacency $\{B,C,D\}$, as shown in Fig. 3(b). This operation brings by-products (dotted circles) which can be further purified by Gaussian operations. We then perform a CV controlled-$Z$ gate between modes $A$ and $D$ to make them disconnected. Therefore, we apply two local shearing operations towards the opposite direction on modes $A$ and $D$ respectively, and then obtain the final four-mode hypergraph state with two 3rd-order hyperedges. In this way, we can directly determine the graphical evolution process of the state for on-demand generation of non-Gaussian quantum state, without complicated calculation.
Furthermore, it is possible to increase the cardinality of the graph by using the presented graphical calculus. In Fig. 4, we show an example where we turn a 3rd-order hypergraph state into a 4th-order hypergraph state. The initial state has two 3rd-order hyperedges $\{A,B,C\}$ and $\{A,D,E\}$ which intersect in mode $A$. By applying a local position shearing operation on mode $A$, we generate a 4th-order hyperedge $\{B,C,D,E\}$. That is, by performing a local Gaussian operation on a 3rd-order hypergraph state, we can obtain a four-mode nonlinearity process which is extremely difficult to achieve in experiments. Note that the current resulting state as shown in Fig. 4(b) is not a standard hypergraph state, but we can remove these by-products in a similar way via auxiliary modes.
Figure 4: Generation of a hypergraph state with a hyperedge of 4th order out of a state with two hyperedges of 3rd order by applying local momentum shearing operation on mode $A$. (a) The initial hypergraph state $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}}\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{q}}}_{D}\operatorname{\mathit{\hat{q}}}_{E}}\ket{0}_{p}^{\otimes 5}$. (b) After performing momentum shearing on mode $A$, the desired blue hyperedge which represents the term $e^{i\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}\operatorname{\mathit{\hat{q}}}_{D}\operatorname{\mathit{\hat{q}}}_{E}}$ of $\{B,C,D,E\}$ is generated. Additionally, two purple dashed lines representing high-order moments, $e^{i\operatorname{\mathit{\hat{q}}}_{B}^{2}\operatorname{\mathit{\hat{q}}}_{C}^{2}/2}$ and $e^{i\operatorname{\mathit{\hat{q}}}_{D}^{2}\operatorname{\mathit{\hat{q}}}_{E}^{2}/2}$, are produced as by-products.
Finally, we demonstrate the possibility of generating the action of the cubic phase gate $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}^{3}}$ [43, 15] on specific states by using similar methods. The cubic phase gate can be decomposed by $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}^{3}}=\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{p}}}_{C}}\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}^{2}\operatorname{\mathit{\hat{q}}}_{C}}\operatorname{e}^{-i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{p}}}_{C}}\operatorname{e}^{-i\operatorname{\mathit{\hat{q}}}_{A}^{2}\operatorname{\mathit{\hat{q}}}_{C}}$, including the term $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}^{2}\operatorname{\mathit{\hat{q}}}_{C}}=\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{p}}}_{B}}\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}}\operatorname{e}^{-i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{p}}}_{B}}\operatorname{e}^{-i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}}$. One term $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}}$ contained therein can be implemented through a CV Toffoli gate introduced by the first example. The other term $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{p}}}_{C}}$ can be achieved by performing shearing operations on modes $B$ and $C$, alongside the control-$Z$ gates executed on modes $A$ and $B$, modes $A$ and $C$, respectively. So, by using the graphical calculus, hypergraph states with Gaussian operations result in a feasible protocol for the implementation of the cubic phase gate, thereby expanding the toolbox for its experimental realization.
Conclusion and Outlook.— We have extended the concept of weighted hypergraph states and developed a graphical calculus tailored for CV non-Gaussian states. This calculus starts from an initial hypergraph state, upon which the action of all single-mode Gaussian operations and Gaussian measurements can be directly described. Then by combining it with two-mode controlled-$Z$ gates as a fundamental building block, we formulated a comprehensive set of rules governing the action of multi-mode Gaussian operations on standard hypergraph states. Our graphical framework offers a visually intuitive way of manipulating CV non-Gaussian states, thereby presenting a valuable toolset that could significantly facilitate the design of quantum circuits in the realm of quantum computation. Accordingly, it is anticipated that this graphical approach can be extended to encompass a broader class of approximate hypergraph states with finite squeezing, which fosters a deeper understanding of the underlying structure and dynamics of non-Gaussian states in CV systems.
Acknowledgments.— We thank Darren W. Moore for discussions. This work was supported by the National Natural Science Foundation of China (No. 12125402, No. 12350006), the Innovation Program for Quantum Science and Technology (No. 2021ZD0301500), Beijing Natural Science Foundation (Grant No. Z240007), the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation, project numbers 447948357 and 440958198), the Sino-German Center for Research Promotion (Project M-0294), the German Ministry of Education and Research (Project QuKuK, BMBF Grant No. 16KIS1618K), and the Stiftung der Deutschen Wirtschaft.
Appendix A Appendix A: Proofs
In this appendix, we provide the derivation details of the results presented in the main text. Since the operations $C_{e}$ commute for all hyperedges $e\in E$, the identity
$\prod_{e\in E}C_{e}(t_{e})=\prod_{e\in E}\operatorname{e}^{it_{e}\operatorname{\mathit{\hat{q}}}_{e}}=\operatorname{e}^{i\sum_{e\in E}t_{e}\operatorname{\mathit{\hat{q}}}_{e}}$ (4)
holds. We can divide the set by a vertex $a\in V$ into disjoint subsets $E=E_{a}\dot{\bigcup}E_{\neg a}$, where $E_{a}=\{e\mid e\in E,a\in e\}$ is the set of hyperedges containing $a$, and $E_{\neg a}=\{e\mid e\in E,a\notin e\}$ is the set of hyperedges not containing $a$. We therefore have
$\prod_{e\in E}C_{e}(t_{e})$ $=\prod_{e\in E_{a}}C_{e}(t_{e})\prod_{e^{\prime}\in E_{\neg a}}C_{e^{\prime}}(t_{e^{\prime}})$
$=\operatorname{e}^{i\sum_{e\in E_{a}}t_{e}\operatorname{\mathit{\hat{q}}}_{e}}\prod_{e^{\prime}\in E_{\neg a}}C_{e^{\prime}}(t_{e^{\prime}})$
$=\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{a}\sum_{e\in\operatorname{\mathcal{A}}(a)}t_{e\cup a}\operatorname{\mathit{\hat{q}}}_{e}}\prod_{e^{\prime}\in E_{\neg a}}C_{e^{\prime}}(t_{e^{\prime}})$
$=\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{a}h(a)}\prod_{e^{\prime}\in E_{\neg a}}C_{e^{\prime}}(t_{e^{\prime}}),$ (5)
where we define $h(a)\coloneqq\sum_{e\in\operatorname{\mathcal{A}}(a)}t_{e\cup a}\operatorname{\mathit{\hat{q}}}_{e}$. Note that by definition $E_{a}=\{e\cup a|e\in\operatorname{\mathcal{A}}(a)\}$.
A.1 Displacement in momentum
The momentum displacement operator $Z^{a}(s)=\operatorname{e}^{is\operatorname{\mathit{\hat{q}}}_{a}}$ modifies the hyperedge set by $(E,t)\rightarrow(E,t)\cup(\{a\},s)$. That is, it adds the hyperedge $\{a\}$ with the weight $s$ if there was no such hyperedge before or changes the weight by $s$.
For $\{a\}\in E$:
$Z^{a}(s)\ket{H}$ $=\operatorname{e}^{is\operatorname{\mathit{\hat{q}}}_{a}}\operatorname{e}^{it_{a}\operatorname{\mathit{\hat{q}}}_{a}}\prod_{e\in E\setminus\{a\}}C_{e}(t_{e})\ket{0}^{\otimes|V|}$
$=\operatorname{e}^{i(s+t_{a})\operatorname{\mathit{\hat{q}}}_{a}}\prod_{e\in E\setminus\{a\}}C_{e}(t_{e})\ket{0}^{\otimes|V|}$
$=C_{a}(s+t_{a})\prod_{e\in E\setminus\{a\}}C_{e}(t_{e})\ket{0}^{\otimes|V|}.$ (6)
The case $\{a\}\notin E$ directly follows by setting $t_{a}=0$ in the above calculation.
A.2 Displacement in position
Applying the position displacement operator $X(s)$ on the mode $a$ modifies the hyperedges in the adjacency $\operatorname{\mathcal{A}}(a)$ of $a$ such that $(E^{\prime},t^{\prime})=(E,t)\cup(\operatorname{\mathcal{A}}(a),\tilde{t})$, where $\tilde{t}_{e}\coloneqq-s\times t_{e\cup a}$. In other words, it adds or updates the hyperedges in the adjacency $\operatorname{\mathcal{A}}(a)$ of $a$ and multiplies the weight of the hyperedge containing $a$ by $s$:
$X^{a}(s)\ket{H}$ $=\operatorname{e}^{-is\operatorname{\mathit{\hat{p}}}_{a}}\prod_{e\in E}C_{e}(t_{e})\ket{0}_{p}^{\otimes\lvert V\rvert}$
$=\operatorname{e}^{-is\operatorname{\mathit{\hat{p}}}_{a}}\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{a}h(a)}\prod_{e^{\prime}\in E_{\neg a}}C_{e^{\prime}}(t_{e^{\prime}})\ket{0}_{p}^{\otimes\lvert V\rvert}$
$=\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{a}h(a)}\operatorname{e}^{-is\operatorname{\mathit{\hat{p}}}_{a}}\operatorname{e}^{-ish(a)}\prod_{e^{\prime}\in E_{\neg a}}C_{e^{\prime}}(t_{e^{\prime}})\ket{0}_{p}^{\otimes\lvert V\rvert}$
$=\prod_{e\in E_{a}}C_{e}(t_{e})\prod_{e^{\prime}\in E_{\neg a}}C_{e^{\prime}}(t_{e^{\prime}})\prod_{e^{\prime\prime}\in\operatorname{\mathcal{A}}(a)}C_{e^{\prime\prime}}(-s\cdot t_{e^{\prime\prime}\cup a})\operatorname{e}^{-it\operatorname{\mathit{\hat{p}}}_{a}}\ket{0}_{p}^{\otimes\lvert V\rvert}$
$=\prod_{e\in E}C_{e}(t_{e})\prod_{e^{\prime\prime}\in\operatorname{\mathcal{A}}(a)}C_{e^{\prime\prime}}(-s\cdot t_{e^{\prime\prime}\cup a})\ket{0}_{p}^{\otimes\lvert V\rvert}.$ (7)
A.3 Shearing
Applying shearing in position is trivial. This expression
$D_{q}^{a}(s)\ket{H}$ $=\operatorname{e}^{i\frac{s}{2}\operatorname{\mathit{\hat{q}}}_{a}^{2}}\prod_{e\in E}C_{e}(t_{e})\ket{0}_{p}^{\otimes\lvert V\rvert}$ (8)
cannot be simplified.
Applying shearing in momentum gives us
$D_{p}^{a}(s)\ket{H}$ $=\operatorname{e}^{i\frac{s}{2}\operatorname{\mathit{\hat{p}}}_{a}^{2}}\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{a}h(a)}\prod_{e^{\prime}\in E_{\neg a}}C_{e^{\prime}}(t_{e^{\prime}})\ket{0}_{p}^{\otimes\lvert V\rvert}$
$=\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{a}h(a)+ish(a)\operatorname{\mathit{\hat{p}}}_{a}}\prod_{e^{\prime}\in E_{\neg a}}C_{e^{\prime}}(t_{e^{\prime}})\ket{0}_{p}^{\otimes\lvert V\rvert}$
$=\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{a}h(a)}\operatorname{e}^{ish(a)\operatorname{\mathit{\hat{p}}}_{a}}\operatorname{e}^{i\frac{s}{2}h(a)^{2}}\prod_{e^{\prime}\in E_{\neg a}}C_{e^{\prime}}(t_{e^{\prime}})\ket{0}_{p}^{\otimes\lvert V\rvert}$
$=\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{a}h(a)}\operatorname{e}^{i\frac{s}{2}h(a)^{2}}\prod_{e^{\prime}\in E_{\neg a}}C_{e^{\prime}}(t_{e^{\prime}})\ket{0}_{p}^{\otimes\lvert V\rvert}$
$=\operatorname{e}^{i\frac{s}{2}h(a)^{2}}\prod_{e\in E}C_{e}(t_{e})\ket{0}_{p}^{\otimes\lvert V\rvert}$
$=\operatorname{e}^{i\frac{s}{2}(\sum_{e\in\operatorname{\mathcal{A}}(a)}t_{e\cup a}\operatorname{\mathit{\hat{q}}}_{e})^{2}}\ket{H}$
$=\prod_{e\in\operatorname{\mathcal{A}}(a)}\operatorname{e}^{i\frac{s}{2}t_{e\cup a}^{2}\operatorname{\mathit{\hat{q}}}_{e}^{2}}\prod_{e,e^{\prime}\in\operatorname{\mathcal{A}}(a),e\neq e^{\prime}}\operatorname{e}^{is\cdot t_{e\cup a}t_{e^{\prime}\cup a}\operatorname{\mathit{\hat{q}}}_{e}\operatorname{\mathit{\hat{q}}}_{e^{\prime}}}\ket{H},$ (9)
where we have used used the identity $\left[\operatorname{\mathit{\hat{p}}}_{a}^{2},\operatorname{\mathit{\hat{q}}}_{a}\right]=-2i\operatorname{\mathit{\hat{p}}}_{a}$.
A.4 One-mode squeezing
Applying a one-mode squeezing operation $S^{a}(s)=\operatorname{e}^{-i\frac{s}{2}(\operatorname{\mathit{\hat{p}}}_{a}\operatorname{\mathit{\hat{q}}}_{a}+\operatorname{\mathit{\hat{q}}}_{a}\operatorname{\mathit{\hat{p}}}_{a})}=\operatorname{e}^{-\frac{s}{2}}\operatorname{e}^{-is\operatorname{\mathit{\hat{q}}}_{a}\operatorname{\mathit{\hat{p}}}_{a}}$ on the mode $a$ changes the weight of all hyperedges containing $a$ by a factor of $\operatorname{e}^{-s}$:
$S^{a}(s)\ket{H}$ $=e^{-\frac{s}{2}}\operatorname{e}^{-is\operatorname{\mathit{\hat{q}}}_{a}\operatorname{\mathit{\hat{p}}}_{a}}\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{a}h(a)}\prod_{e^{\prime}\in E_{\neg a}}C_{e^{\prime}}(t_{e^{\prime}})\ket{0}_{p}^{\otimes\lvert V\rvert}$
$=e^{-\frac{s}{2}}\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{a}h(a)\sum_{n=0}^{\infty}\frac{1}{n!}(-s)^{n}}\operatorname{e}^{-is\operatorname{\mathit{\hat{q}}}_{a}\operatorname{\mathit{\hat{p}}}_{a}}\prod_{e^{\prime}\in E_{\neg a}}C_{e^{\prime}}(t_{e^{\prime}})\ket{0}_{p}^{\otimes\lvert V\rvert}$
$=e^{-\frac{s}{2}}\operatorname{e}^{i\operatorname{e}^{-s}\operatorname{\mathit{\hat{q}}}_{a}h(a)}\prod_{e^{\prime}\in E_{\neg a}}C_{e^{\prime}}(t_{e^{\prime}})\ket{0}_{p}^{\otimes\lvert V\rvert}$
$=e^{-\frac{s}{2}}\prod_{e\in E_{a}}C_{e}(\operatorname{e}^{-s}t_{e})\prod_{e^{\prime}\in E_{\neg a}}C_{e^{\prime}}(t_{e^{\prime}})\ket{0}_{p}^{\otimes\lvert V\rvert}.$ (10)
We have used the Baker-Campbell-Hausdorff identity
$\operatorname{e}^{X}e^{Y}=e^{(Y+[X,Y]+\frac{1}{2!}[X,[X,Y]]+\frac{1}{3!}[X,[X,[X,Y]]]+...)}e^{X},$ (11)
with $X=-is\operatorname{\mathit{\hat{q}}}_{a}\operatorname{\mathit{\hat{p}}}_{a}$ and $Y=i\operatorname{\mathit{\hat{q}}}_{a}h(a)$.
A.5 Rotation
Applying a rotation operator in general gives us a state which is not in the form of a hypergraph state. The general form is given by:
$R^{a}(s)\ket{H}$ $=\operatorname{e}^{i\frac{s}{2}(\operatorname{\mathit{\hat{q}}}_{a}^{2}+\operatorname{\mathit{\hat{p}}}_{a}^{2})}\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{a}h(a)}\prod_{e^{\prime}\in E_{\neg a}}C_{e^{\prime}}(t_{e^{\prime}})\ket{0}_{p}^{\otimes\lvert V\rvert}$
$=\operatorname{e}^{ih(a)(\cos(s)\operatorname{\mathit{\hat{q}}}_{a}+\sin(s)\operatorname{\mathit{\hat{p}}}_{a})}\operatorname{e}^{i\frac{s}{2}(\operatorname{\mathit{\hat{q}}}_{a}^{2}+\operatorname{\mathit{\hat{p}}}_{a}^{2})}\prod_{e^{\prime}\in E_{\neg a}}C_{e^{\prime}}(t_{e^{\prime}})\ket{0}_{p}^{\otimes\lvert V\rvert}$
$=\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{a}h(a)\cos(s)}\operatorname{e}^{i\operatorname{\mathit{\hat{p}}}_{a}h(a)\sin(s)}\operatorname{e}^{i\frac{1}{4}h(a)^{2}\sin(2s)}\prod_{e^{\prime}\in E_{\neg a}}C_{e^{\prime}}(t_{e^{\prime}})\operatorname{e}^{i\frac{s}{2}(\operatorname{\mathit{\hat{q}}}_{a}^{2}+\operatorname{\mathit{\hat{p}}}_{a}^{2})}\ket{0}_{p}^{\otimes\lvert V\rvert}$
$=\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{a}h(a)\cos(s)}\operatorname{e}^{i\operatorname{\mathit{\hat{p}}}_{a}h(a)\sin(s)}\operatorname{e}^{i\frac{1}{4}h(a)^{2}\sin(2s)}\prod_{e^{\prime}\in E_{\neg a}}C_{e^{\prime}}(t_{e^{\prime}})\ket{0}_{p}^{\otimes\lvert V\rvert-1}(R^{a}(s)\ket{0}_{p,a}),$ (12)
where $R^{a}(s)\ket{0}_{p}=-\sin(s)\ket{0}_{q}+\cos(s)\ket{0}_{p}$. We used the Baker-Campbell-Hausdorff identity shown in Eq. 11 with $X=i\frac{s}{2}(\operatorname{\mathit{\hat{q}}}^{2}+\operatorname{\mathit{\hat{p}}}^{2})$ and $Y=i\operatorname{\mathit{\hat{q}}}h(a)$.
Another special case is the Fourier transform $F=R(\frac{\pi}{2})$, which swaps $\operatorname{\mathit{\hat{q}}}$ with $\operatorname{\mathit{\hat{p}}}$, and $\operatorname{\mathit{\hat{p}}}$ with $-\operatorname{\mathit{\hat{q}}}$.
A.6 Controlled-$Z$ gate
The effect of application of the controlled-$Z$ $C_{a,b}(s)=\operatorname{e}^{is\operatorname{\mathit{\hat{q}}}_{a}\operatorname{\mathit{\hat{q}}}_{b}}$ gate on two modes $a,b$ directly follows from the definition of hypergraph states. If the edge $\{a,b\}\in E$ was part of the edge set, its weight gets modified, since
$C_{\{a,b\}}(s)C_{\{a,b\}}(t_{\{a,b\}})=C_{\{a,b\}}(s+t_{\{a,b\}})$ (13)
holds. Otherwise, if $\{a,b\}\notin E$ it adds the edge $\{a,b\}$ with weight $s$ to the edge set.
A.7 Measurement in $\operatorname{\mathit{\hat{q}}}$ basis
Measuring |$\outerproduct{m}{m}_{q,a}$ disconnects mode $a$ from the hypergraph by disconnecting hyperedges containing $a$. The hyperedge set and weights get updated as $E^{\prime}=(E\setminus E_{a})\cup\operatorname{\mathcal{A}}(a)$ and for $e\in E^{\prime}$
$t^{\prime}_{e}=\begin{cases}mt_{e\cup a}&\text{if }e\in\operatorname{\mathcal{A}}(a),\\ t_{e}&\text{else}.\end{cases}$ (14)
In detail we have
[…]
=|$=\outerproduct{m}{m}_{q,a}\operatorname{e}^{imh(a)}\prod_{e^{\prime}\in E_{\neg a}}C_{e^{\prime}}(t_{e^{\prime}})\ket{0}_{p}^{\otimes\lvert V\rvert}$
$=\prod_{e\in\operatorname{\mathcal{A}}(a)}C_{e}(mt_{e\cup a})\prod_{e^{\prime}\in E_{\neg a}}C_{e^{\prime}}(t_{e^{\prime}})\outerproduct{m}{m}_{q,a}\ket{0}_{p}^{\otimes\lvert V\rvert}$
=|$=\ket{m}_{q,a}\prod_{e\in\operatorname{\mathcal{A}}(a)}C_{e}(mt_{e\cup a})\prod_{e^{\prime}\in E_{\neg a}}C_{e^{\prime}}(t_{e^{\prime}})\ket{0}_{p}^{\otimes\lvert V\rvert-1}.$ (15)
A.8 Measurement in $\operatorname{\mathit{\hat{p}}}$ basis
Measuring |$\outerproduct{m}{m}_{p,a}$ disconnects mode $a$ from the hypergraph by disconnecting hyperedges containing $a$ and creates several hyperedges in a superposition of all possible weights, with a phase determined by the weight and the measurement result $m$. The hyperedge set and weight set get updated as $E^{\prime}=(E\setminus E_{a})\cup\operatorname{\mathcal{A}}(a)$ and for $e\in E^{\prime}$
$t^{\prime}(e)=\begin{cases}xt(e\cup a)&\text{if }e\in\operatorname{\mathcal{A}}(a),\\ t(e)&\text{else}.\end{cases}$ (16)
In detail we have
[…]
The integral cannot be further simplified, so it is not a hypergraph state after measuring $\operatorname{\mathit{\hat{p}}}$.
Figure 5: Generation of a state with $m$ hyperedges of order 3 out of a state with one hyperedge of order 3. (a) The initial hypergraph state. (b) After performing momentum shearing on mode $A$, the desired purple hyperedges of $\{B,C,D_{1}\}$,…,$\{B,C,D_{m}\}$ are generated. Additionally, $\operatorname{e}^{i\frac{1}{2}\operatorname{\mathit{\hat{q}}}_{D_{1}}^{2}}$,…, $\operatorname{e}^{i\frac{1}{2}\operatorname{\mathit{\hat{q}}}_{D_{m}}^{2}}$, and $e^{i\frac{1}{2}\operatorname{\mathit{\hat{q}}}_{B}^{2}\operatorname{\mathit{\hat{q}}}_{C}^{2}}$, are produced as by-products. (c) Removing the $\{A,D_{1}\}$, …, $\{A,D_{m}\}$ edge by performing a series of two-mode controlled-$Z$ gates, $C_{A,D_{1}}(-1)$, …, $C_{A,D_{m}}(-1)$. (d) After performing the shearing operations towards the opposite direction on modes $A$ and $D_{1}$, …, $D_{m}$ respectively, the new hypergraph state $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}}\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}\operatorname{\mathit{\hat{q}}}_{D_{1}}}$, …, $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}\operatorname{\mathit{\hat{q}}}_{D_{m}}}\ket{0}_{p}^{\otimes{m+3}}$ is obtained.
Appendix B Appendix B: Further example
In this section we give a further extension of the example given in the main text: In the main text in Figure 3, we couple mode $D$ with mode $A$ and use the graphical rules to show that it generates a new hyperedge of order 3 between modes $B$, $C$, and $D$. An extended example is shown in Fig. 5. We couple multiple modes $D_{1}$, …, $D_{m}$ with mode $A$ and perform the same sequence of operations which has the same effect than applying $m$ Toffoli gates on the initial state. We can see that, after performing the shearing operations on mode $A$ and a series of controlled-$Z$ gates as well as the shearing operations towards the opposite direction on modes $A$ and $D_{1}$, …, $D_{m}$ respectively, the new hypergraph state Instead of applying a non-Gaussian Toffoli gate on a certain hypergraph state, we can apply a sequence of Gaussian operations on the same state which has the same effect. In the main text in Figure 3, we couple mode $D$ with mode $A$ and use the graphical rules to show that it generates a new hyperedge of order 3 between modes $B$, $C$, and $D$. An extended example is shown in Fig. 5. We couple multiple modes $D_{1}$, …, $D_{m}$ with mode $A$ and perform the same sequence of operations which has the same effect than applying $m$ Toffoli gates on the initial state. We can see that, after performing a series of shearing operations and controlled-$Z$ gates, the new hypergraph state $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}}\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}\operatorname{\mathit{\hat{q}}}_{D_{1}}},\dots,\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}\operatorname{\mathit{\hat{q}}}_{D_{m}}}\ket{0}_{p}^{\otimes{m+3}}$ is obtained.